OpenStudy (anonymous):

Find all solutions to the equation. sin x = sqrt(3)/2

OpenStudy (anonymous):

x = pi/3 + k(2)pi AND x = 2pi/3+ k(2)pi also written as x = 60 + k360 AND x = 120 + k360 because "y", which is sin x, is positive in both quadrant I AND II

OpenStudy (anonymous):

The "+" within each of the answers comes from the cyclic nature of the trig functions. "k" is any integer and represents the number of times "around the circle".

OpenStudy (anonymous):

So would the answer be the first one (pictured in the attachment)

OpenStudy (anonymous):

Actually, it's the 3rd one. The key is to look for the pi/3 and the 2pi/3.

OpenStudy (anonymous):

ahh. I see what you mean now. Thanks.

OpenStudy (anonymous):

You're getting this!